### v1

# Kummer-Artin-Schreier-Witt Theory

Abstract
We prove that the Artin-Schreier-Witt isogeny in characteristic $p>0$ lifts to characteristic 00.
14H30, 14H10, 11S15
keywords:
Ramification theory, non-archimedean geometry, Kummer-Artin-Schreier-Witt theory.
1 Introduction
Our motivation is the lifting problem for Galois covers of curves, which asks: given a finite group $G$ and a $G$-cover $\phi$ of a curve in characteristic $p$, does $\phi$ lift to characteristic 00? It is important to note that, in general, the answer is NO. As with many problems in Galois theory, it is natural to begin by investigating the case where $G$ is abelian. There have been many breakthroughs recently; for instance, Obus, Wewers, and Pop have proved that all cyclic covers lift [8, 9].
Recall that every abelian cover can be realized as the pullback of a suitable isogeny of algebraic groups [12]. Therefore, to study the lifting problem for abelian covers, one can instead explore the lift of cyclic isogenies of group schemes from characteristic $p$ to characteristic 00. A classical argument allows us to further restrict to the case where $G$ is cyclic of order $p^{s}$ and the isogeny Weierstrass-$\wp$ of the ring $W_{s}$ of Witt vectors of length $s$, which gives the following exact sequence
$0\xrightarrow{}\mathbb{Z}/p^{s}\xrightarrow{}W_{s}\xrightarrow{\wp}W_{s}\rightarrow 0$ (1)
of group schemes over $\overline{k}$, where ${\rm char}(k)=p$. Sekiguchi and Suwa showed that (1) lifts .
Theorem 1.1 ((c.f. [11, Theorem 8.1])).
For each positive integer $s$, there exists a flat group scheme $\mathcal{W}_{s}$ over $R:=\mathbb{Z}_{(p)}[\zeta_{p^{s}}]$ that fits into the exact sequence
$0\rightarrow\mathbb{Z}/p^{s}\rightarrow\mathcal{W}_{s}\rightarrow\mathcal{V}_{s}:=\mathcal{W}_{s}/(\mathbb{Z}/p^{s})\rightarrow 0$ (2)
of group schemes. The generic fiber of (2) is isomorphic to the Kummer exact sequence
$0\xrightarrow{}\mu_{p^{s}}\xrightarrow{}\mathbb{G}_{m}\xrightarrow{(\cdot)^{p^{s}}}\mathbb{G}_{m}\xrightarrow[]{}0,$ (3)
over $\overline{K}$, where $K=\operatorname{Frac}(R)$. Its special fiber is isomorphic to the Artin-Schreier-Witt sequence (1) over $\overline{\mathbb{F}}_{p}$. Furthermore, $\mathcal{W}_{s}$, as a scheme, is the spectrum of
$R\bigg{[}Y_{0},\ldots,Y_{s-1},\frac{1}{1+\lambda Y_{0}},\frac{1}{E_{2}(Y_{0})+\lambda Y_{1}},\ldots,\frac{1}{E_{s}(Y_{0},\ldots,Y_{s-2})+\lambda Y_{s-1}}\bigg{]},$ (4)
for explicitly determined polynomials $E_{2},\ldots,E_{s}$ and $\lambda:=\zeta_{p}-1$. Also, $\mathcal{V}_{s}$ is the spectrum of
$R\bigg{[}X_{0},\ldots,X_{s-1},\frac{1}{1+p\lambda X_{0}},\frac{1}{G_{2}(X_{0})+p\lambda X_{1}},\ldots,\frac{1}{G_{s}(X_{0},\ldots,X_{s-2})+p\lambda X_{s-1}}\bigg{]},$ (5)
for explicitly determined polynomials $G_{2},\ldots,G_{s}$.
One immediate implication is the following result.
Theorem 1.2 ((c.f. [11, Theorem 3.8])).
If $B$ is a flat local $R$-algebra, then any unramified $\mathbb{Z}/p^{s}$-cover $\operatorname{Spec}C\rightarrow\operatorname{Spec}B$ is given by a Cartesian diagram of the form
${\operatorname{Spec}C}$${\mathcal{W}_{s}}$${\operatorname{Spec}B}$${\mathcal{V}_{s}}$
However, the paper by Sekiguchi and Suwa remains unpublished. The preprint is 90 pages long and quite challenging for us to read (see [6] for a summary instead). Around the same time, Matsuda developed his own version of Kummer-Artin-Schreier-Witt theory from the perspective of $p$-adic differential equations [4]. More precisely, let $\mathbb{K}$ be a henselian local field of mixed characteristic with imperfect residue field $\kappa$. Matsuda established that the canonical restriction morphism $\operatorname{H}^{1}(\kappa,\mathbb{Z}/p^{s})\rightarrow\operatorname{H}^{1}(\mathbb{K},\mathbb{Z}/p^{s})$ is induced by the morphism
$\begin{split}W_{s}(\mathcal{R}^{\operatorname{int}})&\xrightarrow[]{\Psi_{s}}(\mathcal{R}^{\operatorname{int}})^{\times},\\ \underline{X}&\mapsto E_{s,p}(p^{s}\underline{X}),\end{split}$ (6)
where $\mathcal{R}^{\operatorname{int}}$ is a henselian DVR over $\mathbb{K}$, and $E_{s,p}((X_{0},\ldots,X_{s-1}))\in\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket X_{0},\ldots,X_{s-1}\rrbracket$ is a deformation of the Artin-Hasse exponential function.
Specifically, the morphism (6) allows one to study a cyclic extension in mixed characteristic defined by a Kummer class, such as with $E_{s,p}(p^{s}\underline{X})$, by examining the associated Witt vector $\underline{X}$. This approach often simplifies the analysis, as evidenced by many results on Galois extensions in equal-characteristic $p$ obtained through the study of Witt vectors [5, 1].
In this manuscript, we prove the above results using the robust framework developed by Matsuda. More precisely, to recover Theorem 1.1, we apply an algebraization process to (6) to obtain a polynomial $E_{s}(\underline{X})\in\mathbb{Z}_{(p)}[\zeta_{p^{s}}][X_{0},\ldots,X_{s-1}]$, which is a suitable truncation of $E_{s,p}(p^{s}\underline{X})$. We then show that, for $R=\mathbb{Z}_{(p)}[\zeta_{p^{s}}]$, the cover
$Z_{s}^{p^{s}}=E_{s}((X_{0},\ldots,X_{s-1}))$
of $\mathbb{P}^{s}_{R}=\operatorname{Proj}R[X_{0},\ldots,X_{s-1},Y]$ gives (2) after removing the ramification divisor.
Remark 1.3.
Our polynomials $E_{i}$ and $G_{i}$ in (4) and (5) are different from those used by Sekiguchi and Suwa in their work. In fact, these polynomials are not unique, as they are approximations of power series arising from Matsuda theory within a certain degree of accuracy.
1.1 Outline
In Section 2, we recall Kummer theory and Artin-Schreier-Witt theory, which classify tamely ramified and wildly ramified cyclic extensions, respectively. In the same section, we restate Matsuda’s theory, which distinguishes our method from the one used by Sekiguchi and Suwa. Section 3 defines the cover on the level of field extension. More precisely, the section constructs the Kummer class of the extension, which is basically the $G_{i}$’s from (5) by truncating certain power series that arise from Matsuda’s theory. Section 4 derives the $E_{i}$’s from (4) using the Kummer extension previously defined. Section 5 shows that our previous constructions give a cyclic étale cover of an affine scheme with goood reduction. Section 6 further proves that the cover is an isogeny. Finally, Section 7 gives the proof of Theorem 1.2.
1.2 Notations
The letter $K$ will always denote a field of mixed characteristic $(0,p)$ with uniformizer $\pi$ and a discrete valuation $\nu:K^{\times}\rightarrow\mathbb{Q}$, normalized so that $\nu(p)=1$. The ring of integers of $K$ will be denoted by $R$, and its residue field $k=R/(\pi)$ is perfect. An example to keep in mind is $K=\mathbb{Q}_{p}[\zeta_{p^{s}}]$, with $\pi=\zeta_{p^{s}}-1$, which has valuation $\nu(\pi)=\frac{1}{p^{s-1}(p-1)}$. Set $\lambda=\zeta_{p}-1$, a constant that will appear frequently.
We fix an algebraic closure $\overline{K}$ of $K$. The symbol $\mathbb{K}$ usually denotes a Laurent series over $K$. The valuation $\nu$ extends to the Gauss valuation on $\mathbb{K}$. For instance, if $\mathbb{K}=K((X))$ and $f(X)=\sum_{i\geq n}a_{i}X^{i}\in\mathbb{K}$, then $\nu(f)=\inf_{i\geq n}\{\nu(a_{i})\}$.
1.3 Acknowledgements
We thank Mézard and Tossici for useful conversations. The first author thanks the National Center for Theoretical Sciences, National Taiwan University for the excellent working conditions.
2 Background
2.1 Artin-Schreier-Witt theory
Suppose $k$ is a field of characteristic $p$. Then we have the short exact sequence:
$\begin{split}0\xrightarrow{}\mathbb{Z}/p^{s}\cong W_{s}(\mathbb{F}_{p})\xrightarrow{}W_{s}&\xrightarrow{\wp}W_{s}\rightarrow 0,\\ \underline{x}&\mapsto F\underline{x}-\underline{x}\end{split}$ (7)
of group schemes over $\overline{k}$, which induces the identification
$\operatorname{H}^{1}(k,\mathbb{Z}/p^{s})\cong W_{s}(k)/\wp(W_{s}(k)).$
One may hence view $\wp:W_{s}\rightarrow W_{s}$ in (7) as a $\mathbb{Z}/p^{s}$-extension of $\mathcal{K}:=\overline{k}(x_{0},\ldots,x_{s-1})$ given by the Witt vector $(x_{0},\ldots,x_{s-1})$. Furthermore, if Weierstrass-$\wp$ is associated with a character $\overline{\chi}\in\operatorname{H}^{1}(\mathcal{K},\mathbb{Z}/p^{s})$, then $\overline{\chi}^{\otimes p^{s-j}}$ is the $\mathbb{Z}/p^{j}$-subextension corresponding to the Witt vector $(x_{0},\ldots,x_{j-1})$.
2.2 Kummer Theory
Fix $m\in\mathbb{Z}_{\geq 0}$. Suppose $K$ contains a primitive $m$-th root of unity and has characteristic prime to $m$. By Kummer theory, we have the following isomorphism:
$\operatorname{H}^{1}(K,\mathbb{Z}/m)\cong K^{\times}/(K^{\times})^{m}.$
The isomorphism follows from the exact sequence over $\overline{K}$:
$0\rightarrow\mu_{m}\rightarrow\overline{K}^{\times}\xrightarrow{(\cdot)^{m}}\overline{K}^{\times}\rightarrow 0$
Thus, each $\chi\in\operatorname{H}^{1}(K,\mathbb{Z}/p^{s}\mathbb{Z})$ can be associated with a rational function $G$ in $K^{\times}/(K^{\times})^{p^{s}}$. Additionally, the function $G$ can be factored as follows:
$G:=\prod_{i=0}^{s-1}G_{i}^{p^{i}},$
where each $G_{i}\in K\setminus K^{p}$ or $G_{i}=1$. As in Section 2.1, $\chi^{\otimes p^{s-j}}$ corresponds to a character in $\operatorname{H}^{1}(K,\mathbb{Z}/p^{j}\mathbb{Z})$ defined by the Kummer extension:
$Z_{j-1}^{p^{j}}=\prod_{i=0}^{j-1}G_{i}^{p^{i}}.$
In fact, $\chi$ can be defined by a system of Kummer extensions of degree $p$:
$Z_{i}^{p}=Z_{i-1}G_{i},\quad i=0,\ldots,s-1,$
where $Z_{-1}=1$.
We thus establish the following isomorphism:
$(K^{\times})^{\oplus s}/\phi_{s}((K^{\times})^{\oplus s})$ ${\sim}K^{\times}/(K^{\times})^{p^{s}}\xrightarrow[]{\sim}\operatorname{H}^{1}(K,\mathbb{Z}/p^{s}\mathbb{Z})$
$(a_{1},\ldots,a_{s})$ $\mapsto[a_{1}a_{2}^{p}\ldots a_{s}^{p^{s-1}}]\mapsto\mathfrak{K}_{s}(a_{1}a_{2}^{p}\ldots a_{s}^{p^{s-1}}),$
where $\phi_{s}(a_{1},\ldots,a_{s})=(a_{1}^{p},a_{2}^{p}a_{1}^{-1},\ldots,a_{s}^{p}a_{s-1}^{-1})$.
Remark 2.1.
$(K^{\times})^{\oplus s}/\phi_{s}((K^{\times})^{\oplus s})$ is a group under coordinate-wise multiplication. We say $(a_{1},\ldots,a_{s})\sim(b_{1},\ldots,b_{s})$ if there exists $(c_{1},\ldots,c_{s})$ such that
$(a_{1},\ldots,a_{s})=(b_{1},\ldots,b_{s})\cdot(c_{1}^{p},c_{2}^{p}c_{1}^{-1},\ldots,c_{s}^{p}c_{s-1}^{-1}).$
This equivalence relation is induced from $K^{\times}/(K^{\times})^{p^{s}}$.
This leads to a Kummer-like sequence:
$0\xrightarrow{}\mu_{p^{s}}\xrightarrow{}(\mathbb{G}_{m})^{s}\xrightarrow{\phi_{s}}(\mathbb{G}_{m})^{s}\xrightarrow[]{}0,$ (8)
where the injection sends $1$ to $(\zeta_{p},\zeta_{p^{2}},\ldots,\zeta_{p^{s}})$.
2.3 Kummer-Artin-Schreier-Witt theory
In this section, we introduce Matsuda’s theory as presented in [3].
2.3.1 Deformed Artin-Hasse exponential function
We first introduce a deformation of the Artin-Hasse exponential function and some of its properties.
Definition 2.2.
The Artin-Hasse exponential series at $p$ is the formal power series
$E_{p}(t):=\exp\left(\sum_{i=0}^{\infty}\frac{t^{p^{i}}}{p^{i}}\right)\in\mathbb{Z}_{(p)}\llbracket t\rrbracket.$
Define
$E_{s,p}(t):=\frac{E_{p}(\zeta_{p^{s}}t)}{E_{p}(t)}=\exp\left(\sum_{i=0}^{s-1}(\zeta_{p^{s-i}}-1)\frac{t^{p^{i}}}{p^{i}}\right)\in\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket t\rrbracket$ (9)
and, for $\underline{a}=(a_{0},\ldots,a_{s-1})$,
$E_{s,p}(\underline{a})=\prod_{i=0}^{s-1}E_{s-i,p}(a_{i})\in\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket a_{0},\ldots,a_{s-1}\rrbracket.$ (10)
One shows that
$E_{s,p}(\underline{a})E_{s,p}(\underline{b})=E_{s,p}(\underline{a}+\underline{b}).$ (11)
Definition 2.3.
We define the formal power series
$G_{s,p}(\underline{a})=\frac{E_{s,p}(p\underline{a})}{E_{s-1,p}(\underline{a})}\in\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket a_{0},\dots,a_{s-1}\rrbracket.$
We have
$G_{s,p}(\underline{a})G_{s,p}(\underline{b})=G_{s,p}(\underline{a}+\underline{b}).$ (12)
Lemma 2.4.
We have $G_{s,p}(\underline{a})\in\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket(\zeta_{p^{s}}-1)a_{0},\ldots,\lambda a_{s-1}\rrbracket$.
The series can be expressed in the below forms:
$G_{s,p}(\underline{a})$ $=\prod_{j=0}^{s-1}\exp\left(\sum_{i=0}^{s-j-1}(p(\zeta_{p^{s-j-i}}-1)-(\zeta_{p^{s-j-1-i}}-1))\frac{a_{j}^{p^{i}}}{p^{i}}\right)$
$=\prod_{j=0}^{s-1}\exp\left(\sum_{i=0}^{s-j-1}(\zeta_{p^{s-j-i}}-1)(p-1-\zeta_{p^{s-j-i}}-\cdots-\zeta_{p^{s-j-i}}^{p-1})\frac{a_{j}^{p^{i}}}{p^{i}}\right)$
$=\prod_{j=0}^{s-1}E_{s-j,p}\left((p-1-[\zeta_{p^{s-j}}]-\cdots-[\zeta_{p^{s-j}}^{p-1}])[a_{j}]\right),$
where $[a]=(a,0,\ldots,0)\in W_{s}(\mathbb{Z}_{(p)}[\zeta_{p^{s}}])$ is the Teichmüller lift of $a$.
Remark 2.5.
The $E_{i}$ (resp. $G_{i}$) from Theorem 1.1 are approximations of $E_{i,p}$ (resp. $G_{i,p}$).
2.3.2 Matsuda’s Kummer-Artin-Schreier-Witt theory
Let $K$ be a discrete valuation field containing $\zeta_{p^{s}}$, with residue field $k$. The Robba ring $\mathcal{R}$ over $K$ is defined as the ring of Laurent series that converge in some annulus $\{z\mid\rho<\lvert z\rvert<1\}$ for some $10<\rho<$ 1. Let $\mathcal{R}^{\rm bd}$ be the subring of $\mathcal{R}$ consisting of formal sums with bounded coefficients. The ring $\mathcal{R}^{\rm bd}$ carries a Gauss norm:
$\left\lvert\sum_{i\in\mathbb{Z}}a_{i}z^{i}\right\rvert=\max_{i}\{\lvert a_{i}\rvert\}.$
Let $\mathcal{R}^{\operatorname{int}}$ be the subring of $\mathcal{R}^{\rm bd}$ consisting of elements with Gauss norm at most $1$.
Lemma 2.6 ([4, Proposition 3.2]).
The ring $\mathcal{R}^{\operatorname{int}}$ is a Henselian discrete valuation ring. Consequently, $\mathcal{R}^{\rm bd}$ is a Henselian local field with residue field $\kappa:=k((z))$.
Therefore, there is a canonical restriction map
$\operatorname{H}^{1}(\kappa,\mathbb{Z}/p^{s})\rightarrow\operatorname{H}^{1}(\mathcal{R}^{\rm bd},\mathbb{Z}/p^{s}).$ (13)
Theorem 2.7 ([4, Theorem 3.8]).
The homomorphism in (13) is induced by a homomorphism
$W_{s}(\mathcal{R}^{\operatorname{int}})$ ${\Psi_{s}}(\mathcal{R}^{\operatorname{int}})^{\times},$ (14)
$\underline{a}$ $\mapsto E_{s,p}(p^{n}\underline{a}).$
The morphism (14) effectively establishes a unification between Kummer and Artin-Schreier-Witt theories by relating a Witt vector $\underline{a}$ with a Kummer class $E_{s,p}(p^{n}\underline{a})$. Our strategy in this paper is to define the cover using Kummer theory while using an associated Witt vector to prove that the cover satisfies the desired conditions.
Remark 2.8.
We believe Matsuda’s theory has potential applications far beyond the scope of this paper. For instance, expanding the domain and the range of (14) enables us to address inseparable extensions. This will be explored in our future work on the refined Swan conductor [2], an important invariant that measures how far a wildly ramified cover is from having separable reduction.
3 Constructing the Kummer extension
3.1 The Witt Vector Associated to a Kummer Extension
Given a $\mathbb{Z}/p^{s}$-cover of $\mathbb{P}^{s}_{R}=\operatorname{Proj}R[X_{0},\ldots,X_{s-1},Y]$ described by the Kummer equation
$Z_{s}^{p^{s}}=\prod_{i=0}^{s-1}(u_{i})^{p^{i}}=:u,$ (15)
where $\mathbb{K}$ is the function field of $\mathbb{P}^{s}_{R}$ and $u_{i}\in(\mathcal{R}^{\operatorname{int}})^{\times}\subset\mathbb{K}$. Let $\underline{X}^{\prime}=(X_{0}^{\prime},\ldots,X_{s-1}^{\prime})\in W_{s}(\mathcal{R}^{\operatorname{int}})$ be a Witt vector such that $\Psi_{s}(\underline{X}^{\prime})=u$, where $\Psi_{s}$ is Matsuda’s morphism (14). In this section, we prove the existence of the $u_{i}$ such that $\underline{X}^{\prime}\equiv(X_{0},\ldots,X_{s-1})\pmod{\pi}$.
The proof of the following result is straightforward from the definition of $E_{i,p}$.
Proposition 3.1.
Suppose that $m<s$. Then
$E_{s,p}((0,\ldots,0,X_{0}^{\prime},\ldots,X_{m-1}^{\prime}))=E_{s-m,p}((X_{0}^{\prime},\ldots,X_{m-1}^{\prime})).$
Proposition 3.2.
Recall that $\lambda:=\zeta_{p}-1$. With the above setting, we can compute the entries $X_{i}^{\prime}$ of an associated Witt vector inductively as follows:
$X_{0}^{\prime}=\frac{1}{p\lambda}\ln(u_{0}),\;X_{1}^{\prime}=\frac{1}{p\lambda}\ln\left(\frac{u_{1}}{G_{2,p}((X_{0}^{\prime},0))}\right),\;\ldots,\;X_{s-1}^{\prime}=\frac{1}{p\lambda}\ln\left(\frac{u_{s-1}}{G_{s,p}((X_{0}^{\prime},\ldots,X_{s-2}^{\prime},0))}\right).$
Proof.
Let us proceed by induction on $i$. Suppose first that $i=1$. Then it is necessary that $\Psi_{1}(X_{0}^{\prime})=\exp(p\lambda X_{0}^{\prime})=u_{0}$. Therefore, $X_{0}^{\prime}=\frac{1}{p\lambda}\ln(u_{0})$ as expected.
Assume the induction hypothesis for $s-1$. That is, one can find $X_{0}^{\prime},\ldots,X_{s-2}^{\prime}$ such that $E_{s-1,p}(p^{s-1}(X_{0}^{\prime},\ldots,X_{s-2}^{\prime}))=\prod_{i=0}^{s-2}(u_{i})^{p^{i}}$. We want to find $X_{s-1}^{\prime}$ such that
$\begin{split}\Psi_{s}((X_{0}^{\prime},\ldots,X_{s-1}^{\prime}))=E_{s,p}(p^{s}(X_{0}^{\prime},\ldots,X_{s-1}^{\prime}))&=\prod_{i=0}^{s-1}(u_{i})^{p^{i}}\\ &=E_{s-1,p}(p^{s-1}(X_{0}^{\prime},\ldots,X_{s-2}^{\prime}))(u_{s-1})^{p^{s-1}}.\end{split}$
Since we only need to find a solution, we take the $p^{s-1}$-th root of both sides, which gives
$\begin{split}E_{s,p}(p(X_{0}^{\prime},\ldots,X_{s-1}^{\prime}))&=E_{s-1,p}(X_{0}^{\prime},\ldots,X_{s-2}^{\prime})u_{s-1},\\ E_{s,p}(p(X_{0}^{\prime},\ldots,X_{s-2}^{\prime},0))E_{s,p}(p(0,\ldots,0,X_{s-1}^{\prime}))&=E_{s-1,p}(X_{0}^{\prime},\ldots,X_{s-2}^{\prime})u_{s-1}.\end{split}$
The second equality follows from the decomposition of Witt vectors
$\underline{x}=(x_{0},\ldots,x_{s-1},x_{s},x_{s+1},\ldots)=(x_{0},\ldots,x_{s-1},0,0,\ldots)+(\underbrace{0,\ldots,0}_{s\text{ times}},x_{s},x_{s+1},\ldots)$
and the commutativity of $E_{s,p}$ by (11). Due to Proposition 3.1, we have
$E_{s,p}(p(0,\ldots,0,X_{s-1}^{\prime}))=E_{1,p}(p(X_{s-1}^{\prime})).$
Therefore,
$\exp(p\lambda X_{s-1}^{\prime})=u_{s-1}\frac{E_{s-1,p}((X_{0}^{\prime},\ldots,X_{s-2}^{\prime}))}{E_{s,p}(p(X_{0}^{\prime},\ldots,X_{s-2}^{\prime},0))}=\frac{u_{s-1}}{G_{s,p}((X_{0}^{\prime},\ldots,X_{s-2}^{\prime},0))}.$
Thus,
$X_{s-1}^{\prime}=\frac{1}{p\lambda}\ln\left(\frac{u_{s-1}}{G_{s,p}((X_{0}^{\prime},\ldots,X_{s-2}^{\prime},0))}\right),$
which completes the induction. ∎
3.2 Construct the Kummer Extension
Let us now further assume that
$u_{i}:=G_{i+1}(X_{0},\ldots,X_{i-1})+p\lambda X_{i},$ (16)
for some $G_{i+1}(X_{0},\ldots,X_{i-1})\in\mathbb{Z}_{(p)}[\zeta_{p^{i+1}}][X_{0},\ldots,X_{i-1}]$. Set
$G^{\prime}_{i+1}:=G_{i+1,p}((X_{0}^{\prime},\ldots,X_{i-1}^{\prime},0))-G_{i+1}(X_{0},\ldots,X_{i-1}),$
where $X_{0}^{\prime},\ldots,X_{i-1}^{\prime}$ are derived from $u_{0},\ldots u_{i-1}$ and Proposition 3.2.
Proposition 3.3.
In the above settings, suppose $\nu(G_{i+1}^{\prime})>\frac{p}{p-1}$ for $i=0,\ldots,s$. Then $X_{i}^{\prime}\equiv X_{i}\mod(\zeta_{p^{i+1}}-1)$ for all $i$.
Proof.
The case $i=0$ is straightforward. Suppose the proposition holds for $i<s$. Applying Proposition 3.2, we obtain
$X_{s-1}^{\prime}$ $=\frac{1}{p\lambda}\ln\left(\frac{G_{s}(X_{0},\ldots,X_{s-2})+p\lambda X_{s-1}}{G_{s,p}((X_{0}^{\prime},\ldots,X_{s-2}^{\prime},0))}\right)$
$=\frac{1}{p\lambda}\ln\left(1+\frac{-G^{\prime}_{s}+p\lambda X_{s-1}}{G_{s,p}((X_{0}^{\prime},\ldots,X_{s-2}^{\prime},0))}\right)$
$=\frac{1}{p\lambda}\ln\left(1-G^{\prime}_{s}+p\lambda X_{s-1}+o\left(p^{p/(p-1)}\right)\right)$
$=X_{s-1}+\frac{1}{p\lambda}o\left(p^{p/(p-1)}\right),$
where $o\left(p^{p/(p-1)}\right)$ refers to terms with valuation strictly larger than $\frac{p}{p-1}$. The third equality follows from the fact that $\nu(G_{s,p}((X_{0}^{\prime},\ldots,X_{s-2}^{\prime},0))-1)\geq 1/(p^{s}(p-1))$ as a series in $\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket X_{0}^{\prime},\allowbreak\ldots,X_{s-2}^{\prime}\rrbracket$ due to Lemma 2.4 and $X_{j}\equiv X^{\prime}_{j}\pmod{(\zeta_{p^{j+1}}-1)}$ for $j<s-1$ by induction, which allows us to write
$\frac{1}{G_{s,p}((X_{0}^{\prime},\ldots,X_{s-2}^{\prime},0))}=1+O\left(p^{1/(p^{s}(p-1))}\right)$
as a series in $\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket X_{0},\allowbreak\ldots,X_{s-2}\rrbracket$. Here $O\left(p^{1/(p^{s}(p-1))}\right)$ refers to terms with valuation at least $1/(p^{s}(p-1))$. In addition, since $\nu(-G^{\prime}_{s}+p\lambda X_{s-1}+o\left(p^{p/(p-1)}\right))\geq\frac{p}{p-1}$, we have, for $m>1$,
$\nu\left(\frac{\left(-G^{\prime}_{s}+p\lambda X_{s-1}+o\left(p^{p/(p-1)}\right)\right)^{m}}{m}\right)>\frac{p}{p-1}.$
The final equality then follows from the series expansion of the $\ln(1+X)$ function. ∎
Proposition 3.3 suggests that one may construct the polynomial $G_{i+1}$ in (16) as an approximation of $G_{i+1,p}((X_{0}^{\prime},\ldots,X_{i-1}^{\prime},0))$.
Definition 3.4.
Let
$f=\sum_{n_{0},\ldots,n_{s}\geq 0}a_{n_{0},\ldots,n_{s}}X_{0}^{n_{0}}\cdots X_{s}^{n_{s}}\in R\llbracket X_{0},\ldots,X_{s}\rrbracket$
satisfy the Tate condition if, for every multi-index $n=(n_{0},\ldots,n_{s})$, the coefficients $a_{n}$ satisfy:
$\nu(a_{n})\to\infty\quad\text{as}\quad|n|=n_{0}+\cdots+n_{s}\to\infty.$
Proposition 3.5.
For each $i$, one can inductively compute $G_{i+1}(X_{0},\ldots,X_{i-1})\in\mathbb{Z}_{(p)}[\zeta_{p^{i+1}}]\allowbreak[X_{0},\ldots,X_{i-1}]$ such that:
  1. 1.
    $G_{i+1,p}((X_{0}^{\prime},\ldots,X_{i-1}^{\prime},0))\in\mathbb{Z}_{(p)}[\zeta_{p^{i+1}}]\llbracket X_{0},\ldots,X_{i}\rrbracket$ satisfies the Tate condition,
  2. 2.
    $\nu(G_{i+1}^{\prime})>\frac{p}{p-1}$ and $G_{i+1}^{\prime}\in\mathbb{Z}_{(p)}[\zeta_{p^{i+1}}]\llbracket X_{0},\ldots,X_{i}\rrbracket$ satisfies the Tate condition, and
  3. 3.
    $X_{i}^{\prime}$ satisfies the Tate condition as a series in $\mathbb{Z}_{(p)}[\zeta_{p^{i+1}}]\llbracket X_{0},\ldots,X_{i}\rrbracket$.
Proof.
Let us first consider the case $i=1$. Then we can simply pick $u_{0}=1+p\lambda X_{0}$ or $G_{1}=1$. By Proposition 3.2, the Witt vector associated with the order $p$ Kummer class $u_{0}$ is $(X_{0}^{\prime})$ where
$\begin{split}X_{0}^{\prime}&=\frac{1}{p\lambda}\ln(1+p\lambda X_{0})\\ &=\frac{1}{p\lambda}\left(p\lambda X_{0}-\frac{p^{2}\lambda^{2}X_{0}^{2}}{2}+\cdots+(-1)^{j+1}\frac{p^{j}\lambda^{j}X_{0}^{j}}{j}+\cdots\right)\end{split}$
The valuations of the coefficients are given by
$\nu\left(\frac{p^{j}\lambda^{j}}{p\lambda j}\right)=\begin{cases}j-1+\frac{j-1}{p-1}=\frac{p(j-1)}{p-1},&p\nmid j\\ p^{i}l-1+\frac{p^{i}l-1}{p-1}-i=\frac{p(p^{i}l-1)}{p-1}-i,&j=p^{i}l,\;(p,l)=1\end{cases}$
Hence, it is clear that the coefficient of $X_{0}^{j}$ has valuation greater than 00 when $j>1$, and $X_{0}^{\prime}$ satisfies the Tate condition. This completes the base case.
Now, consider the case $i=s$. By Lemma 2.4, $G_{s,p}((X_{0}^{\prime},\ldots,X_{s-2}^{\prime},0))$, regarded as a power series in $\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket X_{0}^{\prime},\ldots,X_{s-2}^{\prime}\rrbracket$, satisfies the Tate condition. Furthermore, by the induction hypothesis, $X_{j}^{\prime}$ satisfies the same condition for $j=0,\ldots,s-2$ as a series over $\mathbb{Z}_{(p)}[\zeta_{p^{j+1}}]\llbracket X_{0},\ldots,X_{j}\rrbracket$. Therefore, $G_{s,p}((X_{0}^{\prime},\ldots,X_{s-2}^{\prime},0))$ satisfies the condition as a series in $\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket X_{0},\ldots,X_{s-2}\rrbracket$, proving 1.
Next, we define $G_{s}\in\mathbb{Z}_{(p)}[\zeta_{p^{s}}][X_{0},\ldots,X_{s-2}]$ to be the polynomial obtained by truncating $G_{s,p}((X_{0}^{\prime},\ldots,X_{s-2}^{\prime},0))$ at terms with valuations at most $\frac{p}{p-1}$. Assertion 2 is immediately verified. Assertion 3 can be derived from the computation of $X_{s-1}^{\prime}$ in the proof of Proposition 3.3. Recall that
$X_{s-1}^{\prime}$ $=\frac{1}{p\lambda}\ln\left(1+\frac{-G^{\prime}_{s}+p\lambda X_{s-1}}{G_{s,p}((X_{0}^{\prime},\ldots,X_{s-2}^{\prime},0))}\right)=:\frac{1}{p\lambda}\ln(1+U).$
It is clear that both the numerator and the denominator of $U$ satisfy the condition, so $U$ itself does as well. The claim that $\ln(1+U)$ verifies the condition follows immediately. ∎
3.3 Defining the étale Cover
Definition 3.6.
Define $R:=\mathbb{Z}_{(p)}[\zeta_{p^{s}}]$ and set $\pi:=\zeta_{p^{s}}-1$. Let $\mathcal{V}_{s}$ be an open subset of the projective space $\mathbb{P}^{s}_{R}=\operatorname{Proj}R[X_{0},\ldots,X_{s-1},Y]$ given by
$\operatorname{Spec}R\left[X_{0},\ldots,X_{s-1},\frac{1}{1+p\lambda X_{0}},\frac{1}{G_{2}(X_{0})+p\lambda X_{1}},\ldots,\frac{1}{G_{s}(X_{0},\ldots,X_{s-2})+p\lambda X_{s-1}}\right],$
where each $G_{i}$ is a polynomial constructed in Proposition 3.5. Let $\mathcal{W}_{s}$ be the integral closure of $\mathcal{V}_{s}$ over the extension of $K(X_{0},\ldots,X_{s-1})$ defined by
$Z_{s-1}^{p^{s}}=\prod_{i=0}^{s-1}(G_{i+1}(X_{0},\ldots,X_{i-1})+p\lambda X_{i})^{p^{i}},$ (17)
where $G_{1}=1$.
Due to Theorem 2.7 and Proposition 3.3, the special fiber of $\mathcal{W}_{s}\rightarrow\mathcal{V}_{s}$ is birational to the Artin-Schreier-Witt extension of $\operatorname{Proj}k[x_{0},\ldots,x_{s-1},y]$, where $x_{i}$ (resp. $y$) is the reduction of $X_{i}$ (resp. $Y$) modulo $\pi$, given by the Witt vector $(x_{0},\ldots,x_{s-1})$.
The following result follows directly from Kummer theory.
Proposition 3.7.
Specifically, the cover of $\mathbb{P}^{s}_{R}$ defined by (17) is branched precisely along the hypersurfaces given by $G_{i+1}(X_{0},\ldots,X_{i-1})+p\lambda X_{i}=0$ for $10\leq i\leq s-$ 1. Consequently, the cover $\mathcal{W}_{s}\rightarrow\mathcal{V}_{s}$ is unramified.
4 Compute the $\mathcal{W}_{s}$
In this section, we analyze the structure of the $\mathbb{Z}/p^{s}$-cover $\mathcal{W}_{s}\rightarrow\mathcal{V}_{s}$ defined in the previous section. Specifically, we compute the polynomials $E_{i}$ in (4).
4.1 The Case $s=1$
As before, $\mathcal{W}_{1}\rightarrow\mathcal{V}_{1}$ is given by
$Z_{0}^{p}=1+p\lambda X_{0}.$ (18)
Consider the change of variables $Z_{0}=1+\lambda Y_{0}$. This substitution transforms (18) into
$\begin{split}1+p\lambda X_{0}&=(1+\lambda Y_{0})^{p}\\ &=1+\lambda^{p}Y_{0}^{p}+p\lambda Y_{0}+\sum_{j=2}^{p-1}\binom{p}{j}\lambda^{j}Y_{0}^{j}.\end{split}$
Hence, it simplifies to
$Y_{0}+\frac{\lambda^{p-1}}{p}Y_{0}^{p}+o\left(p^{0}\right)=X_{0}.$ (19)
Here, $o\left(p^{0}\right)$ denotes terms with valuation strictly larger than zero. Note that, since
$\frac{\lambda^{p-1}}{p}\equiv-1\pmod{\lambda}$
[13, Section 2.4], the equation above reduces modulo $\lambda$ to the Artin-Schreier equation
$y_{0}-y_{0}^{p}=x_{0}.$
To generalize this change of variables for the case $s>1$, we follow the strategy of associating a Witt vector as in the previous section. Recall that the right-hand side of (18), due to Proposition 3.3, is associated with the Witt vector $(X_{0}^{\prime})$, where $X_{0}^{\prime}\in\mathbb{Z}_{(p)}[\zeta_{p}]\llbracket X_{0}\rrbracket$ is congruent to $X_{0}$ modulo $\lambda$. We can rewrite the defining equation as
$Z_{0}^{p}=E_{1,p}(p(X_{0}^{\prime})).$ (20)
Note that if we can write $Z_{0}=E_{1,p}(Y_{0}^{\prime})$, then, due to (27), $Y_{0}^{\prime}=X_{0}^{\prime}$. Therefore, it makes sense to define the change of variables as an approximation of
$E_{1,p}(Y_{0}^{\prime})=\exp(\lambda Y_{0}^{\prime})=1+\lambda Y_{0}^{\prime}+\frac{(\lambda Y_{0}^{\prime})^{2}}{2!}+\frac{(\lambda Y_{0}^{\prime})^{3}}{3!}+\cdots$
Thus, we choose $Z_{0}=1+\lambda Y_{0}$. Then, as $Z_{0}=E_{1,p}(Y_{0}^{\prime})=\exp(\lambda Y_{0}^{\prime})$, one can write $Y_{0}^{\prime}$ in terms of $Y_{0}$ as follows:
$\begin{split}Y_{0}^{\prime}&=\frac{1}{\lambda}\ln(1+\lambda Y_{0})\\ &=Y_{0}-\frac{\lambda Y_{0}^{2}}{2}+\cdots+(-1)^{p-1}\frac{\lambda^{p-1}Y_{0}^{p}}{p}+\cdots\end{split}$ (21)
It follows that
$Y_{0}^{\prime}\equiv Y_{0}-Y_{0}^{p}\pmod{\lambda}.$
Thus, we can also conclude that $Y_{0}-Y_{0}^{p}-X_{0}\in\lambda\mathbb{Z}_{(p)}[\zeta_{p}]\llbracket X_{0},Y_{0}\rrbracket$.
4.2 The Case $s>1$
In general, for $i=1,\ldots,s$, we aim to construct a system of change of variables
$Z_{i-1}=E_{i}(Y_{0},\ldots,Y_{i-2})+\lambda Y_{i-1},$ (22)
to better understand the structure of $\mathcal{W}_{s}$ and to derive an Artin-Schreier-type equation similar to (19). As in Section 4.1, beginning with the system of $\mathbb{Z}/p$-Kummer extensions
$Z_{i-1}^{p}=Z_{i-2}(G_{i}(X_{0},\ldots,X_{i-2})+p\lambda X_{i-2}),$
we seek a Witt vector $(Y_{0}^{\prime},\ldots,Y_{s-1}^{\prime})$ such that, for $1\leq i\leq s$,
$Z_{i-1}=E_{i,p}((X_{0}^{\prime},\ldots,X_{i-1}^{\prime})).$
Proposition 4.1.
With the above notations, the components $Y_{i}^{\prime}$ of the Witt vector $(Y_{0}^{\prime},\ldots,Y_{s-1}^{\prime})$ can be computed inductively as follows:
$Y_{0}^{\prime}$ $=\frac{1}{\lambda}\ln\left(1+\lambda Y_{0}\right),$
$Y_{1}^{\prime}$ $=\frac{1}{\lambda}\ln\left(\frac{E_{2}(Y_{0})+\lambda Y_{1}}{E_{2,p}((Y_{0}^{\prime},0))}\right),$
$\vdots$
$Y_{s-1}^{\prime}$ $=\frac{1}{\lambda}\ln\left(\frac{E_{s}(Y_{0},\ldots,Y_{s-2})+\lambda Y_{s-1}}{E_{s,p}((Y_{0}^{\prime},\ldots,Y_{s-2}^{\prime},0))}\right).$
Proof.
By definition, we have
$\begin{split}E_{s}(Y_{0},\ldots,Y_{s-2})+\lambda Y_{s-1}&=E_{s,p}((\underline{Y}_{s}^{\prime}))\\ &=E_{s,p}((Y_{0}^{\prime},\ldots,Y_{s-2}^{\prime},0))\cdot E_{s,p}((0,\ldots,0,Y_{s-1}^{\prime}))\\ &=E_{s,p}((Y_{0}^{\prime},\ldots,Y_{s-2}^{\prime},0))\cdot E_{1,p}((Y_{s-1}^{\prime}))\\ &=E_{s,p}((Y_{0}^{\prime},\ldots,Y_{s-2}^{\prime},0))\cdot\exp\left(\lambda Y_{s-1}^{\prime}\right).\end{split}$
The desired formula for $Y_{s-1}^{\prime}$ follows immediately. ∎
The formula suggests that each $E_{i}$ in (22) should be defined in terms of $E_{i,p}((Y_{0}^{\prime},\ldots,Y_{i-2}^{\prime},0))$.
Proposition 4.2.
For each $i$, there exists a polynomial $E_{i}(Y_{0},\ldots,Y_{i-2})\in\mathbb{Z}_{(p)}[\zeta_{p^{i}}][Y_{0},\ldots,Y_{i-2}]$ such that, applying the change of variables (22), the associated Witt vector $(Y_{0}^{\prime},\ldots,Y_{s-1}^{\prime})$ satisfies the relation:
$Y_{i-1}^{\prime}=(-1)^{p}\frac{\lambda^{p-1}}{p}Y_{i-1}^{p}+Y_{i-1}+F_{i}(Y_{0},\ldots,Y_{i-1}),$
where $F_{i}\in\mathbb{Z}_{(p)}[\zeta_{p^{i}}]\llbracket Y_{0},\ldots,Y_{i-1}\rrbracket$, and $F_{i}$ reduces to a polynomial in $\mathbb{F}_{p}[y_{0},\ldots,y_{i-2}]$ modulo $\zeta_{p^{i}}-1$.
The proposition holds for $i=1$. We assume it holds for $i\leq s-1$.
4.3 Approximation of the Power Series $E_{s,p}$
Proposition 4.3.
All terms of $E_{i,p}((X_{0},\ldots,X_{i-1}))$, aside from the constant term $1$, have positive valuations that are sharply bounded below by $\frac{1}{p^{i}(p-1)}$.
Proof.
We begin by computing the valuations for the case $i=1$. Then
$\begin{split}E_{1,p}((X_{0}))&=\exp(\lambda X_{0})\\ &=1+\sum_{i\geq 1}\frac{\lambda^{i}X_{0}^{i}}{i!}.\end{split}$
It is well-known that $\nu(n!)=\sum_{j\geq 1}\left\lfloor\frac{n}{p^{j}}\right\rfloor=\frac{n-s_{p}(n)}{p-1}$, where $s_{p}(n)$ denotes the sum of digits of $n$ in base $p$. Therefore, the valuations of the coefficients are given by
$\nu\left(\frac{\lambda^{i}}{i!}\right)=\begin{cases}\frac{i}{p-1}-\frac{i-s_{p}(i)}{p-1}=\frac{s_{p}(i)}{p-1}\geq\frac{1}{p-1},&p\nmid i,\\ \frac{p^{n}}{p-1}-\frac{p^{n}-1}{p-1}=\frac{1}{p-1},&i=p^{n}.\end{cases}$
This proves the base case, so we have $\nu(E_{1,p}((X_{0}))-1)=\frac{1}{p-1}$.
For $i>1$, using the commutativity of $E_{i,p}$, we find
$\begin{split}E_{i,p}(p(X_{0},\ldots,X_{i-1}))&=E_{i,p}((pX_{0},0,\ldots,0))\cdot E_{i,p}((0,\tilde{X}_{1},\ldots,\tilde{X}_{i-1}))\\ &=E_{i,p}((pX_{0},0,\ldots,0))\cdot E_{i-1,p}((\tilde{X}_{1},\ldots,\tilde{X}_{i-1})),\end{split}$
where $\tilde{X}_{j}=X_{j-1}^{p}+pR$. By the induction hypothesis, $\nu(E_{i-1,p}((\tilde{X}_{1},\ldots,\tilde{X}_{i-1}))-1)=\frac{1}{p^{i-2}(p-1)}$, which is strictly less than 1 in general, except when $p=2$ and $i=2$. Thus, $\nu(E_{i,p}(p(X_{0},\ldots,X_{i-1}))-1)=\frac{1}{p^{i-2}(p-1)}$ for $p>2$ or $p=2$ and $i>2$. A direct computation confirms that $\nu(E_{2,2}((X_{0},X_{1}))-1)=\frac{1}{2}$.
Now, suppose $E_{i,p}((X_{0},\ldots,X_{i-1}))=1+A$, where $A\in R:=\mathbb{Z}_{(p)}[\zeta_{p^{i}}]\llbracket X_{0},\ldots,X_{i-1}\rrbracket$. Then
$E_{i,p}(p(X_{0},\ldots,X_{i-1}))-1=A^{p}+\sum_{j=1}^{p-1}\binom{p}{j}A^{j}.$
Since $\nu(E_{i,p}(p(X_{0},\ldots,X_{i-1}))-1)=\frac{1}{p^{i-2}(p-1)}<1$ unless $p=2$ and $i=2$, we deduce that $\nu(A^{p}+\sum_{j=1}^{p-1}\binom{p}{j}A^{j})=p\nu(A)$. Therefore, we conclude that $\nu(E_{i,p}((X_{0},\ldots,X_{i-1}))-1)=\frac{1}{p^{i-1}(p-1)}$, completing the proof. ∎
Proposition 4.4.
Suppose $f(X)=1+\sum_{j\geq 0}a_{j}X^{j}\in R\llbracket X\rrbracket$, with $\nu(a_{j})>0$ and the valuations bounded below by a number $B$. Suppose moreover that $f(X)^{p}$ has a finite number of terms whose coefficients have valuations strictly smaller than $p/(p-1)$. Then $f(X)$ itself has only finitely many terms whose coefficients have valuations strictly smaller than $1/(p-1)$.
Proof.
We can write
$\left(\sum_{j\geq 0}a_{j}X^{j}\right)^{p}=\sum_{n\geq 0}\left(\sum_{\begin{subarray}{c}m_{1}+m_{2}+\dots+m_{r}=p\\ m_{1}i_{1}+m_{2}i_{2}+\dots+m_{r}i_{r}=n\end{subarray}}\frac{p!}{m_{1}!m_{2}!\dots m_{r}!}\cdot a_{i_{1}}^{m_{1}}a_{i_{2}}^{m_{2}}\cdots a_{i_{r}}^{m_{r}}\right)X^{n}=:\sum_{n\geq 0}c_{n}X^{n}.$
To compute the valuation of each summand in the inner sum, we find
$\nu\left(\frac{p!}{\prod_{k}m_{k}!}\cdot a_{i_{1}}^{m_{1}}a_{i_{2}}^{m_{2}}\cdots a_{i_{r}}^{m_{r}}\right)=1+\sum_{j=1}^{r}m_{j}\nu(a_{i_{j}})-\sum_{k=1}^{r}\nu(m_{k}!).$
Consider the cases based on $r$:
  1. 1.
    If $r=1$: Then $m_{1}=p$ and $n=pi_{1}$. Thus, the valuation is $p\nu(a_{i_{1}})$.
  2. 2.
    If $r>1$: This case arises when $n$ is not divisible by $p$. Then $\nu(m_{k}!)=0$, so the valuation becomes $1+\sum_{j=1}^{r}m_{j}\nu(a_{i_{j}})$.
Assume there exists an index $m>\frac{l}{p}$ such that $\nu(a_{m})<\frac{1}{p-1}$. Then $a_{m}^{p}$ contributes to $c_{pm}$. For $\nu(c_{pm})\geq\frac{p}{p-1}$, there must be distinct indices $i_{1},\ldots,i_{s}<pm$ such that $\sum_{j=1}^{r^{\prime}}m_{j}i_{j}=pm$, where $\sum_{j=1}^{r^{\prime}}m_{j}=p$, which implies that
$1+\sum_{j=1}^{r^{\prime}}m_{j}\nu(a_{i_{j}})\leq p\nu(a_{m}),$ (23)
which is impossible if $\nu(a_{m})\leq\frac{1}{p}$. Thus, we conclude that $\nu(a_{m})>\frac{1}{p}$ for $m>\frac{l}{p}$.
Additionally, due to the assumption $\nu(a_{m})<\frac{1}{p-1}$, we have the relation
$p\nu(a_{m})-1-\nu(a_{m})=(p-1)\nu(a_{m})-1<1-1=0,$
which implies $p\nu(a_{m})-1<\nu(a_{m})$.
Furthermore, for $\sum_{j=1}^{r^{\prime}}m_{j}i_{j}=pm>l$, at least one $i_{s}>\frac{l}{p}$ for some $s$. Therefore, $\nu(a_{i_{s}})>\frac{1}{p}$. Finally, as the lower bound of the valuations is $B$, it follows from (23) that
$\nu(a_{i_{s}})\leq p\nu(a_{m})-1-(p-1)B<\nu(a_{m})-(p-1)B.$
This leads to coefficients with progressively smaller valuations until reaching a value not exceeding $\frac{1}{p}$. Therefore, we can conclude that $\nu(a_{m})\geq\frac{1}{p-1}$ for all $m>\frac{l}{p}$, completing the proof. ∎
Proposition 4.5.
Proposition 4.4 extends to the case $f\in R\llbracket X_{0},\ldots,X_{i-1}\rrbracket$, where $i>1$.
Proof.
We may view $f$ as a series in $R\llbracket X_{0},\ldots,X_{i-2}\rrbracket\llbracket X_{i-1}\rrbracket=:R_{i-1}\llbracket X_{i-1}\rrbracket$, where $R_{i-1}$ is equipped with the Gauss valuation. Applying Proposition 4.4 to $f$ in terms of $X_{i-1}$ then allows us to assume that
$f=1+\sum_{j=1}^{n}a_{j}X_{i-1}^{j}+O\left(p^{1/(p-1)}\right),$
where $a_{j}\in R_{i-1}$, and $\nu(a_{j})<1/(p-1)$ for $j=1,\ldots,n$. Furthermore, we have
$f^{p}=(1+\sum_{j=1}^{n}a_{j}X_{i-1}^{j})^{p}+O\left(p^{p/(p-1)}\right).$ (24)
Therefore, $f_{i-1}^{p}:=(1+\sum_{j=1}^{n}a_{j}X_{i-1}^{j})^{p}$ has finitely many terms with valuations smaller than $p/(p-1)$. We then regard $f_{i-1}$ as a series in $R\llbracket X_{0},\ldots,X_{i-3},X_{i-1}\rrbracket\llbracket X_{i-2}\rrbracket=:R_{i-2}\llbracket X_{i-2}\rrbracket$. Applying Proposition 4.4 to $f_{i-1}$ in terms of $X_{i-2}$, we find that
$f_{i-1}=1+\sum_{j=1}^{m}b_{j}X_{i-2}^{j}+O\left(p^{1/(p-1)}\right),$
where $b_{j}\in R\llbracket X_{0},\ldots,X_{i-3}\rrbracket[X_{i-1}]$. Repeating this process $i-2$ times completes the proof. ∎
Proposition 4.6.
$E_{s,p}(p(Y_{0}^{\prime},\ldots,Y_{s-2}^{\prime},0))$ contains only finitely many terms with coefficients of valuation strictly smaller than $p/(p-1)$.
Proof.
By definition, we have
$\begin{split}E_{s,p}(p(Y_{0}^{\prime},\ldots,Y_{s-2}^{\prime},0))&=E_{s-1,p}((Y_{0}^{\prime},\ldots,Y_{s-2}^{\prime}))\cdot G_{s,p}((Y_{0}^{\prime},\ldots,Y_{s-2}^{\prime},0))\\ &=(E_{s-1}(Y_{0},\ldots,Y_{s-3})+\lambda Y_{s-2})\cdot G_{s,p}((Y_{0}^{\prime},\ldots,Y_{s-2}^{\prime},0)).\end{split}$
The first factor meets the condition because it is a polynomial, and the second factor does so by Lemma 2.4 and the fact that $Y_{s}^{\prime}$ satisfies the conditions of Proposition 4.2 by the induction hypothesis. The proposition then follows. ∎
Corollary 4.7.
The series $E_{s,p}((Y_{0}^{\prime},\ldots,Y_{s-2}^{\prime},0))\in R\llbracket Y_{0},\ldots,Y_{s-2}\rrbracket$ contains only finitely many terms with coefficients of valuation strictly smaller than $\frac{1}{p-1}$.
Proof.
The result follows from Proposition 4.3, Proposition 4.5, and Proposition 4.6. ∎
Remark 4.8.
This result does not generally hold for $E_{s,p}((Y_{0},\ldots,Y_{s-2},0))$.
Therefore, the following definition is well-defined.
Definition 4.9.
Define $E_{s}(Y_{0},\ldots,Y_{s-2})\in\mathbb{Z}_{(p)}[\zeta_{p^{s}}][Y_{0},\ldots,Y_{s-2}]$ as the polynomial obtained by truncating $E_{s,p}((Y_{0}^{\prime},\ldots,Y_{s-2}^{\prime},0))\in\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket Y_{0},\ldots,Y_{s-2}\rrbracket$ at terms whose coefficients have valuations strictly smaller than $\frac{1}{p-1}$.
Proposition 4.10.
With the choice of $E_{s}(Y_{0},\ldots,Y_{s-2})$ as in Definition 4.9, the entry $Y_{s-1}^{\prime}$ of the associated Witt vector has the form
$Y_{s-1}^{\prime}=(-1)^{p}\frac{\lambda^{p-1}}{p}Y_{s-1}^{p}+Y_{s-1}+F_{s}(Y_{0},\ldots,Y_{s-1}),$
where $F_{s}(Y_{0},\ldots,Y_{s-1})\in\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket Y_{0},\ldots,Y_{s-1}\rrbracket$.
Proof.
The case $s=1$ is proved in Section 4.1. For $s>1$, we have the following expansion from Proposition 4.1 for $Y_{s-1}^{\prime}$:
$\begin{split}Y_{s-1}^{\prime}&=\frac{1}{\lambda}\ln\left(\frac{E_{s}(Y_{0},\ldots,Y_{s-2})+\lambda Y_{s-1}}{E_{s,p}((Y_{0}^{\prime},\ldots,Y_{s-2}^{\prime},0))}\right)\\ &=\frac{1}{\lambda}\ln\left(1+\frac{S_{s}+\lambda Y_{s-1}}{E_{s,p}((Y_{0}^{\prime},\ldots,Y_{s-2}^{\prime},0))}\right)\\ &=\frac{1}{\lambda}\ln\left(1+\lambda Y_{s-1}+S_{s}+o\left(p^{1/(p-1)}\right)\right)\\ &=Y_{s-1}+(-1)^{p-1}\frac{\lambda^{p-1}}{p}Y_{s-1}^{p}+\frac{S_{s}}{\lambda}+(-1)^{p-1}\frac{S_{s}^{p}}{p\lambda}+o\left(p^{0}\right)\end{split}$
Here, $S_{s}:=E_{s}(Y_{0},\ldots,Y_{s-2})-E_{s,p}((Y_{0}^{\prime},\ldots,Y_{s-2}^{\prime},0))$, which has valuation at least $\frac{1}{p-1}$ by definition. The rest follows similarly to the proof of Proposition 3.3 and the case for $s=1$. This completes the proof. ∎
Proposition 4.11.
The minimal polynomial of $Y_{s-1}$ over $\mathbb{Q}[\zeta_{p^{s}}](Y_{0},\ldots,Y_{s-2},X_{s-1})$ is
$\frac{\lambda^{p-1}}{p}Y_{s-1}^{p}+Y_{s-1}+g_{s-1}(Y_{0},\ldots,Y_{s-1})-(E_{s-1}(Y_{0},\ldots,Y_{s-3})+\lambda Y_{s-2})X_{s-1}=0.$ (25)
In particular, $X_{j}$ lies in $\mathbb{Z}_{(p)}[\zeta_{p^{j+1}}]\left[Y_{0},\ldots,Y_{j},\frac{1}{1+\lambda Y_{0}},\ldots,\frac{1}{E_{j-1}+\lambda Y_{j-1}}\right]$ for $j=0,\ldots,s-1$.
Proof.
Substituting $Z_{s-1}=E_{s}(Y_{0},\ldots,Y_{s-2})+\lambda Y_{s-1}$ into
$Z_{s-1}^{p}=Z_{s-2}(G_{s}(X_{0},\ldots,X_{s-2})+p\lambda X_{s-1})$
transforms the equation to
$\begin{split}\frac{1}{p\lambda}\big{(}(E_{s}(Y_{0},\ldots,Y_{s-2})+\lambda Y_{s-1})^{p}&-(E_{s-1}(Y_{0},\ldots,Y_{s-3})+\lambda Y_{s-2})G_{s}(X_{0},\ldots,X_{s-2})\big{)}\\ &=X_{s-1}(E_{s-1}(Y_{0},\ldots,Y_{s-3})+\lambda Y_{s-2}).\end{split}$ (26)
Let us analyze the left-hand side. By construction, we have
$\begin{split}E_{s}(Y_{0},\ldots,Y_{s-2})&=E_{s,p}((Y_{0}^{\prime},\ldots,Y_{s-2}^{\prime},0))+O\left(p^{1/(p-1)}\right),\\ E_{s-1}(Y_{0},\ldots,Y_{s-3})+\lambda Y_{s-2}&=E_{s-1,p}((Y_{0}^{\prime},\ldots,Y_{s-2}^{\prime})),\\ G_{s}(X_{0},\ldots,X_{s-2})&=G_{s,p}((X_{0}^{\prime},\ldots,X_{s-2}^{\prime},0))+o\left(p^{p/(p-1)}\right).\end{split}$
The first two expressions are series in $\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket Y_{0},\ldots,Y_{s-2}\rrbracket$, and the last is a series in $\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket X_{0},\allowbreak\ldots,X_{s-2}\rrbracket$. Additionally, for $j=0,\ldots,s-1$, we have $X_{j}^{\prime}=Y_{j}^{\prime}$, and $X_{j}$ can be expressed as a polynomial in $\mathbb{Z}_{(p)}[\zeta_{p^{j-1}}]\left[Y_{0},\ldots,Y_{j},\frac{1}{1+\lambda Y_{0}},\ldots,\frac{1}{E_{j-1}+\lambda Y_{j-1}}\right]$ by induction. Thus, we obtain
$G_{s}(X_{0},\ldots,X_{s-2})=G_{s,p}((Y_{0}^{\prime},\ldots,Y_{s-2}^{\prime},0))+o\left(p^{p/(p-1)}\right)$
as an element of $\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket Y_{0},\ldots,Y_{s-2}\rrbracket$. Recall also that
$E_{s,p}(p(Y_{0},\ldots,Y_{s-2},Y_{s-1}))=E_{s-1,p}((Y_{0},\ldots,Y_{s-2}))\cdot G_{s,p}((Y_{0},\ldots,Y_{s-2},Y_{s-1})).$
Therefore, the left-hand side of (26) simplifies to
LHS $=\frac{1}{p\lambda}\bigg{(}\left(E_{s,p}((Y_{0}^{\prime},\ldots,Y_{s-2}^{\prime},0))+O\left(p^{1/(p-1)}\right)+\lambda Y_{s-1}\right)^{p}$
$\quad-E_{s-1,p}((Y_{0}^{\prime},\ldots,Y_{s-2}^{\prime}))\left(G_{s,p}((Y_{0}^{\prime},\ldots,Y_{s-2}^{\prime},0))+o\left(p^{p/(p-1)}\right)\right)\bigg{)}$
$=\frac{\lambda^{p-1}}{p}Y_{s-1}^{p}+Y_{s-1}+O\left(p^{1/(p-1)}\right)$
as a series in $\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket Y_{0},\ldots,Y_{s-1}\rrbracket$. This forces the side to be a polynomial in $\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\big{[}Y_{0},\ldots\allowbreak\ldots,Y_{s-1},\frac{1}{1+\lambda Y_{0}},\ldots,\frac{1}{E_{s-2}+\lambda Y_{s-2}}\big{]}$. Furthermore, the left-hand side is a degree-$p$ expression in $Y_{s-1}$, so it is the minimal polynomial of $Y_{s-1}$ over $\mathbb{Q}[\zeta_{p^{s}}](Y_{0},\ldots,Y_{s-2},X_{s-1})$. Finally, the claim that $X_{s-1}$ is an element of $\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\left[Y_{0},\ldots,Y_{j},\frac{1}{1+\lambda Y_{0}},\ldots,\frac{1}{E_{s-1}+\lambda Y_{s-2}}\right]$ follows immediately. ∎
Proposition 4.12.
The series $F_{s}$ from Proposition 4.10 reduces to a polynomial in $\mathbb{F}_{p}[y_{0},\ldots,y_{s-2}]$ modulo $(\zeta_{p^{s}}-1)$.
Proof.
Since $Y_{s}^{\prime}=X_{s}^{\prime}$, there exists a power series $H_{s-1}\in\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket X_{0},\ldots,X_{s-1}\rrbracket$, which is also a unit, such that
$\begin{split}-uY_{s-1}^{p}+Y_{s-1}+F_{s}&=X_{s-1}H_{s-1},\\ H_{s-1}^{-1}(-uY_{s-1}^{p}+Y_{s-1}+F_{s})(E_{s-1}(Y_{0},\ldots,Y_{s-3})+\lambda Y_{s-2})&=X_{s-1}(E_{s-1}(Y_{0},\ldots,Y_{s-3})+\lambda Y_{s-2}),\end{split}$
where $u=\frac{\lambda^{p-1}}{p}(-1)^{p}$. By Proposition 4.11, we can regard both $X_{s-1}$ and $H_{s-1}^{-1}$ as elements of $\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket Y_{0},\ldots,Y_{s-1}\rrbracket$. Thus, $H_{s-1}^{-1}\equiv 1\pmod{(\zeta_{p^{s}}-1)}$. Consequently, $-uY_{s-1}^{p}+Y_{s-1}+F_{s}$ is equivalent to the left-hand side of (26) modulo $(\zeta_{p^{s}}-1)$.
The result then follows from the expansion
$(E_{s}(Y_{0},\ldots,Y_{s-2})+\lambda Y_{s-1})^{p}=\lambda^{p}Y_{s-1}^{p}+p\lambda Y_{s-1}E_{s}(Y_{0},\ldots,Y_{s-2})+E_{s}^{p}+o\left(p^{p/(p-1)}\right),$
and the fact that $E_{s}(Y_{0},\ldots,Y_{s-2})$ reduces to $1$ modulo $(\zeta_{p^{s}}-1)$. ∎
That concludes the proof of Proposition 4.2. ∎
5 Smoothness of the Integral Closure
In this section, we show that $\mathcal{W}_{s}$ has good reduction over $R:=\mathbb{Z}_{(p)}[\zeta_{p^{s}}]$. Set $\pi:=\zeta_{p^{s}}-1$.
5.1 Case $s=1$
The covering $\mathcal{W}_{1}$ is defined by the diagram in Figure 1, showing the relationships between $\mathcal{W}_{1}$, $\mathcal{V}_{1}$, and their function fields.
[…]
Figure 1: Construction of $\mathcal{W}_{1}$
Recall that $L_{1}/K(X_{0})$ is the Kummer extension given by
$Z_{0}^{p}=1+p\lambda X_{0}.$ (27)
That is, the coordinate ring of $\mathcal{W}_{1}$ is the integral closure of that of $\mathcal{V}_{1}$ in $L_{1}$. Recall also that applying the change of variables $Z_{0}=1+\lambda Y_{0}$ to (27) yields
$Z_{0}^{p}=1+\lambda^{p}Y_{0}^{p}+p\lambda Y_{0}+\sum_{i=2}^{p-1}\binom{p}{i}(\lambda Y_{0})^{i}=1+p\lambda X_{0},$ (28)
which simplifies to
$S_{0}(Y_{0}):=Y_{0}+\frac{\lambda^{p-1}}{p}Y_{0}^{p}+\frac{1}{p}\sum_{i=2}^{p-1}\binom{p}{i}\lambda^{i-1}Y_{0}^{i}=X_{0},$ (29)
and therefore reduces modulo $\lambda$ to
$y_{0}-y_{0}^{p}=x_{0}.$
Proposition 5.1.
$R\left[Y_{0},\frac{1}{1-\lambda Y_{0}}\right]$ is the integral closure of $R\left[X_{0},\frac{1}{1+p\lambda X_{0}}\right]$ in $L_{1}$.
Proof.
It is straightforward to show that $A:=R\left[X_{0},\frac{1}{1+p\lambda X_{0}}\right]$ (resp. $B_{1}:=R\left[Y_{0},\frac{1}{1+\lambda Y_{0}}\right]$) is integrally closed in $\operatorname{Frac}A=K(X_{0})$ (resp. $\operatorname{Frac}B_{1}=K(Y_{0})=L_{1}$), hence normal. From (29) and
$\frac{1}{Z_{0}^{p}}=\left(\frac{1}{1+\lambda Y_{0}}\right)^{p}=\frac{1}{1+p\lambda X_{0}},$
we see that $Y_{0}$ and $\frac{1}{1+\lambda Y_{0}}$ are integral over $A$. Hence, $B_{1}$ is contained in the closure of $A$ in $K(Y_{0})$. Additionally, (29) shows that $X_{0}$ can be expressed as a polynomial in $B_{1}$. Therefore, the canonical inclusion $K(X_{0})\hookrightarrow K(Y_{0})$ induces $A\hookrightarrow B_{1}$. Thus, as $B_{1}$ is normal, it contains the integral closure of $A$ in $K(Y_{0})$. Therefore, $B_{1}$ is the integral closure of $A$ in $L_{1}$. ∎
Thus, $\mathcal{W}_{1}=\operatorname{Spec}R\left[Y_{0},\frac{1}{1-\lambda Y_{0}}\right]$ has good reduction $\operatorname{Spec}\mathbb{F}_{p}[y_{0}]$. Additionally, from (27) and Kummer theory, the cover $\mathcal{Y}_{1}\rightarrow\mathbb{P}^{1}_{R}$ branches generically at $X_{0}=\infty$ and $X_{0}=\frac{1}{p\lambda}$. Therefore, $\mathcal{W}_{1}\rightarrow\mathcal{V}_{1}$ is étale with Galois group $\mathbb{Z}/p$, and its special fiber is isomorphic to the Artin-Schreier extension $\operatorname{Spec}\mathbb{F}_{p}[y_{0}]\rightarrow\operatorname{Spec}\mathbb{F}_{p}[x_{0}]$ given by $-y_{0}^{p}+y_{0}=x_{0}$.
5.2 When $s>1$
Consider Figure 2, which extends Figure 1 to show the step-by-step construction of $\mathcal{W}_{s}$.
[…]
Figure 2: Construction of $\mathcal{W}_{s}$
Here we set $\mathbb{P}^{s}_{R}=\operatorname{Proj}R[X_{0},X_{1},\ldots,X_{s-1},V_{0}]$ and
$\mathcal{O}_{\mathcal{V}_{s}}=R\left[X_{0},\ldots,X_{s-1},\frac{1}{1+p\lambda X_{0}},\ldots,\frac{1}{G_{s}(X_{0},\ldots,X_{s-2})+p\lambda X_{s-1}}\right],$
where the polynomials $G_{i-1}(X_{0},\ldots,X_{i-2})$ are computed in Section 3, and $L_{i+1}/L_{i}$ is given by
$Z_{i}^{p}=Z_{i-1}\left(G_{i+1}(X_{0},\ldots,X_{i-1})+p\lambda X_{i}\right).$ (30)
Recall from Proposition 4.11 that the change of variables $Z_{i-1}=E_{i}(Y_{0},\ldots,Y_{i-2})+\lambda Y_{i-1}$, given in Section 4, leads to the relations
$X_{i-1}=S_{i-1}(Y_{0},\ldots,Y_{i-1}),$
where $S_{i-1}$ is an Artin-Schreier-Witt type polynomial similar to (29) in the ring
$R\left[Y_{0},\ldots,Y_{i-1},\frac{1}{1+\lambda Y_{0}},\ldots,\frac{1}{E_{i-1}(Y_{0},\ldots,Y_{i-3})+\lambda Y_{i-2}}\right],$
for $i=1,\ldots,s-1$. Assume the induction hypothesis that $\mathcal{W}_{s-1}=\operatorname{Spec}B_{s-1}$, with $B_{s-1}$ given by
$R\left[Y_{0},\ldots,Y_{s-2},X_{s-1},\frac{1}{1+\lambda Y_{0}},\ldots,\frac{1}{E_{s-1}(Y_{0},\ldots,Y_{s-3})+\lambda Y_{s-2}},\frac{1}{G_{s}(S_{0},\ldots,S_{i-2})+p\lambda X_{s-1}}\right],$
and the $S_{j}(Y_{0},\ldots,Y_{j})$ as defined above.
Proposition 5.2.
The integral closure of $B_{s-1}$ in $L_{s}$ is
$B_{s}=R\left[Y_{0},\ldots,Y_{s-1},\frac{1}{1+\lambda Y_{0}},\ldots,\frac{1}{E_{s}(Y_{0},\ldots,Y_{s-2})+\lambda Y_{s-1}}\right].$
Proof.
We apply the following change of variables:
$Z_{s-1}=:E_{s}(Y_{0},\ldots,Y_{s-2})+\lambda Y_{s-1},$
where $E_{s}(Y_{0},\ldots,Y_{s-2})$ is defined in Section 4, transforming equation (30) when $i=s-1$ into
$(E_{s}(Y_{0},\ldots,Y_{s-2})+\lambda Y_{s-1})^{p}=(E_{s-1}(Y_{0},\ldots,Y_{s-3})+\lambda Y_{s-2})\left(G_{s}(S_{0},\ldots,S_{i-2})+p\lambda X_{s-1}\right),$ (31)
which can also be rewritten as
$\left(\frac{1}{E_{s}(Y_{0},\ldots,Y_{s-2})+\lambda Y_{s-1}}\right)^{p}=\frac{1}{G_{s}(S_{0},\ldots,S_{i-2})+p\lambda X_{s-1}}\cdot\frac{1}{E_{s-1}(Y_{0},\ldots,Y_{s-3})+\lambda Y_{s-2}}.$
In addition, Proposition 4.11 shows that (31) can be simplified to a minimal polynomial for $Y_{s-1}$ over $B_{s-1}$:
$\frac{1}{E_{s-1}(Y_{0},\ldots,Y_{s-3})+\lambda Y_{s-2}}\left(\frac{\lambda^{p-1}}{p}Y_{s-1}^{p}+Y_{s-1}+g_{s-1}(Y_{0},\ldots,Y_{s-1})\right)=X_{s-1}.$
These equations imply that $Y_{s-1}$ and $\frac{1}{E_{s}+\lambda Y_{s-1}}$ are integral over $B_{s-1}$, and that $X_{s-1}$ lies in $B_{s}$. Therefore, by reasoning similar to that in the proof of Proposition 5.1, we conclude that the integral closure of $B_{s-1}$ in $L_{s}$ is $B_{s}$. ∎
Thus, $B_{s}$ is also the integral closure of $\mathcal{O}_{\mathcal{V}_{s}}$ in $L_{s}$, and $\operatorname{Spec}B_{s}=:\mathcal{W}_{s}\rightarrow\mathcal{V}_{s}$ is étale with a Galois group $\mathbb{Z}/p^{s}$ and has good reduction. In particular, the special fiber $\operatorname{Spec}\mathbb{F}_{p}[y_{0},\ldots,y_{s-1}]\rightarrow\operatorname{Spec}\mathbb{F}_{p}[x_{0},\ldots,x_{s-1}]$ corresponds to the Witt vector $(x_{0},\ldots,x_{s-1})$.
6 The Group Structure
From Proposition 4.1 and Proposition 3.2, we have the (not necessarily algebraic) coordinate maps $\gamma:\mathcal{W}_{s}\rightarrow W_{s}$ and $\lambda:\mathcal{V}_{s}\rightarrow W_{s}$, given by
$(Y_{0},\ldots,Y_{s-1})\mapsto(Y_{0}^{\prime},\ldots,Y_{s-1}^{\prime})$ (32)
and
$(X_{0},\ldots,X_{s-1})\mapsto(X_{0}^{\prime},\ldots,X_{s-1}^{\prime}),$ (33)
respectively.
Definition 6.1.
Define a group structure on $\mathcal{W}_{s}$ (and similarly on $\mathcal{V}_{s}$) by the associated Witt vector group structure on $W_{s}$ via $\gamma$ (or via $\lambda$).
Recall also that $X_{i}^{\prime}\equiv X_{i}\pmod{\pi}$. This indicates the natural group structure on the special fibers, which coincides with the group structure on $W_{s}(\mathbb{F}_{p})$.
Proposition 6.2.
We have a commutative diagram of group morphisms
[…]
Proof.
The commutativity of the first square follows from Definition 6.1.
The maps $\tilde{E}_{s,p}$ and $\tilde{\Psi}_{s}$ are defined by
$\begin{split}(Y_{0}^{\prime},\ldots,Y_{s}^{\prime})&\mapsto(E_{1,p}((Y_{0}^{\prime})),\ldots,E_{s+1,p}((Y_{0}^{\prime},\ldots,Y_{s}^{\prime}))),\text{ and}\\ (X_{0}^{\prime},\ldots,X_{s}^{\prime})&\mapsto(G_{1,p}((X_{0}^{\prime})),\ldots,G_{s+1,p}((X_{0}^{\prime},\ldots,X_{s}^{\prime}))),\end{split}$
respectively. Observe that each $E_{j,p}((Y_{0}^{\prime},\ldots,Y_{j-1}^{\prime}))$ (resp. $G_{j,p}((X_{0}^{\prime},\ldots,X_{j-1}^{\prime}))$) is a polynomial in $\mathbb{Z}_{(p)}[\zeta_{p^{j}}][Y_{0},\ldots,Y_{j-1}]$ (resp. in $\mathbb{Z}_{(p)}[\zeta_{p^{j}}][X_{0},\ldots,X_{j-1}]$). The commutativity of the second square follows from the identity
$\left(E_{i+1,p}((Y_{0}^{\prime},\ldots,Y_{i}^{\prime}))\right)^{p}\left(E_{i,p}((Y_{0}^{\prime},\ldots,Y_{i-1}^{\prime}))\right)^{-1}=Z_{i}^{p}Z_{i-1}^{-1}=G_{i+1,p}((X_{0}^{\prime},\ldots,X_{i}^{\prime})).$
Note that the composition $\mathcal{V}_{s}\rightarrow\mathbb{G}_{m}$ is precisely the morphism $\Psi_{s}((X_{0}^{\prime},\ldots,X_{s-1}^{\prime}))$. Additionally, $\tilde{E}_{s,p}\circ\gamma$ and $\tilde{\Psi}_{s}\circ\lambda$ are morphisms of algebraic group schemes, which are isomorphic on the generic fibers. This completes the proof of the commutative diagram’s structure. ∎
This concludes the proof of Theorem 1.1. ∎
7 The Universality
In this section, we will establish the proof of Theorem 1.2. Our methodology follows standard procedures, as outlined in [11, Section 3] and [7, Section 6.4]. To begin, we will fix some notation. We define $\mathcal{W}_{1}:=\operatorname{Spec}R\left[X_{0},\frac{1}{1+\lambda X_{0}}\right]$. Let $R_{\lambda}:=R/\lambda$ and denote $\mathbb{G}_{m,R_{\lambda}}$ as the special fiber of $\mathbb{G}_{m}:=\operatorname{Spec}R\left[X,\frac{1}{X}\right]$. We have the natural closed immersion $\mathbb{G}_{m,R_{\lambda}}\hookrightarrow\mathbb{G}_{m}$ denoted by $\iota$. Regarding the computation of $\mathcal{W}_{1}$, the following exact sequence of sheaves on the étale site of $\operatorname{Spec}A$ is particularly useful (see [10, Lemma 1.1]):
[…]
where $\alpha:x\mapsto 1+\lambda x$ and $\beta:t\mapsto t\mod\lambda$. The following result is derived from [11, Theorem 2.2].
Lemma 7.1.
Suppose $R$ contains the $p^{s}$-th root of unity. If $B$ is a flat local $R$-algebra, then
$\operatorname{H}^{1}(\operatorname{Spec}B,\mathcal{W}_{1})=0.$
Proof.
By considering the long exact sequence associated with (34) relative to $\operatorname{Spec}B$, we obtain the following:
$0\to\mathcal{W}_{1}(B)\to\mathbb{G}_{m}(B)$ $=B^{*}\to\iota_{*}\mathbb{G}_{m,R_{\lambda}}(B)=(B/(\lambda B))^{*}\to$
$\to\operatorname{H}^{1}(\operatorname{Spec}B,\mathcal{W}_{1})\to\operatorname{H}^{1}(\operatorname{Spec}B,\mathbb{G}_{m})\to\cdots$
By Hilbert’s Theorem 90, we know that $\operatorname{H}^{1}(\operatorname{Spec}B,\mathbb{G}_{m})=0$. Given that the map $f:B^{*}\to(B/(\lambda B))^{*}$ is surjective, the exactness of the long exact sequence implies that $\operatorname{H}^{1}(\operatorname{Spec}B,\mathcal{W}_{1})\cong\operatorname{coker}(f)=0$. ∎
We can employ induction by leveraging the following result.
Lemma 7.2.
For any $s>1$, the following sequence of group schemes
$0\to\mathcal{W}_{1}\to\mathcal{W}_{s}\xrightarrow{}\mathcal{W}_{s-1}\to 0$ (35)
is exact and it fits into the following commutative diagram:
00{0}${\mathbb{Z}_{p}/(p)}$${\mathbb{Z}_{p}/(p^{s})}$${\mathbb{Z}_{p}/(p^{s-1})}$00{0}00{0}${\mathcal{W}_{1}}$${\mathcal{W}_{s}}$${\mathcal{W}_{s-1}}$00{0}
Proof.
The morphism $\mathcal{W}_{s}\to\mathcal{W}_{s-1}$ arises from the inclusion of the $R$-algebras that define the schemes, with the kernel given by
$\operatorname{Spec}R\left[Y_{s-1},\frac{1}{1+\lambda Y_{s-1}}\right]\cong\mathcal{W}_{1}.$
Moreover, the Artin-Schreier-Witt theory establishes the desired commutative diagram. ∎
By induction, we obtain the following result.
Proposition 7.3.
Suppose $R$ contains the $p^{s}$-th root of unity. If $B$ is a flat local $R$-algebra, then for every $s\geq 1$, we obtain
$\operatorname{H}^{1}(\operatorname{Spec}B,\mathcal{W}_{s})=0.$
Proof.
Assume the induction hypothesis that $\operatorname{H}^{1}(\operatorname{Spec}B,\mathcal{W}_{s-1})=0$. We claim that $\operatorname{H}^{1}(\operatorname{Spec}B,\mathcal{W}_{s})$ vanishes. Taking the long exact sequence associated with (35) with respect to $\operatorname{Spec}B$ yields:
$0\rightarrow\mathcal{W}_{1}(B)\rightarrow\mathcal{W}_{s}(B)\rightarrow\mathcal{W}_{s-1}(B)\rightarrow\operatorname{H}^{1}(\operatorname{Spec}B,\mathcal{W}_{1})\rightarrow\\ \rightarrow\operatorname{H}^{1}(\operatorname{Spec}B,\mathcal{W}_{s})\rightarrow\operatorname{H}^{1}(\operatorname{Spec}B,\mathcal{W}_{s-1})\rightarrow\cdots$
Since $\operatorname{H}^{1}(\operatorname{Spec}B,\mathcal{W}_{1})=\operatorname{H}^{1}(\operatorname{Spec}B,\mathcal{W}_{s-1})=0$, we get the desired result. ∎
We will now proceed to prove our main theorem, which states that the group scheme $\mathcal{W}_{s}$ over $R$ provides a unified framework for the Kummer-Artin-Schreier-Witt theory.
Theorem 7.4.
Suppose $R$ contains the $p^{s}$-th root of unity. If $B$ is a flat local $R$-algebra, then any unramified $\mathbb{Z}/p^{s}$-cover $\operatorname{Spec}C\rightarrow\operatorname{Spec}B$ is given by a Cartesian diagram of the form:
${\operatorname{Spec}C}$${\mathcal{W}_{s}}$${\operatorname{Spec}B}$${\mathcal{V}_{s}}$$\scriptstyle{\psi_{s}}$
Proof.
Taking the long exact sequence from the Kummer sequence:
$0\rightarrow\mathbb{Z}/p^{s}\rightarrow\mathcal{W}_{s}\rightarrow\mathcal{V}_{s}\rightarrow 0$
yields
$0\displaystyle 0\to\mathcal{V}_{s}(B)/\mathcal{W}_{s}(B)\rightarrow\mathrm{H}^{1}(\operatorname{Spec}B,\mathbb{Z}/p^{s})\rightarrow\mathrm{H}^{1}(\operatorname{Spec}B,\mathcal{W}_{s})=$ 0 .
It implies that $\mathbb{Z}/p^{s}$-torsors over $\operatorname{Spec}B$ are in one-to-one correspondence with elements of $\mathcal{V}_{s}(B)/\mathcal{W}_{s}(B)$. ∎